Two 2.2 cm -diameter disks face each other, 1.9 mm apart. They are charged to ±20nC.

A.

What is the electric field strength between the disks?

B.

A proton is shot from the negative disk toward the positive disk. What launch speed must the proton have to just barely reach the positive disk?

Thanks
Answer #1

(a) E = Q / oA

where A is area of disk = pi * r2

where , r = 2.2 / 2 = 1.1 cm

E = 20e-9 / 8.85e-12 * pi * 0.0112

E = 5.945e6 N/C

-----------------------------------------------------------------

(b) using Newton's second law

F = ma

qE = ma

a = qE / m

a = 1.6e-19 * 5.945e6 / 1.67e-27

a = 5.695e14 m/s2

now, using kinematics

v2 - u2 = 2ad

as proton comes to stop at positive plate v = 0

u = sqrt (2ad)

u = sqrt ( 2 * 5.695e14 * 1.9e-3)

u = 1.47e6 m/s